# Bi-parental Reproduction may Enhance Species Sustainability by Conserving Shared Parental Traits more Faithfully than Mono-parental Reproduction

### v3

@@ Abstract
There are many recognized effects of sexual reproduction and other forms of bi-parental reproduction on the sustainment and evolution of species. They include increased diversity, accelerated adaptation, constrained accumulation of deleterious mutations, and the homogenization of species’ genotype. Nevertheless, many questions remain open regarding the evolution of bi-parental reproduction. Here we explore some of the finer details of the homogenization effect, which we believe deserve deeper analysis and discussion. We present a first-cut mathematical analysis, which suggests that bi-parental reproduction enhances the retention of shared, common properties of the parent generation in their offspring, as compared with mono-parental, clonal reproduction. For illustration, we accompany the mathematical claim and proof with several simulation examples depicting and comparing bi- and mono-parental reproduction of real numbers, images of printed text, and images of flowers. We argue that the retention of shared traits of preceding generations is an intrinsic effect of merging the encodings of parents’ traits, regardless of physical, chemical, biological and social aspects of the reproduction process and of the specific trait at hand. As survival of offspring often depends on their ability to step in and fill voids left by failing, dying or absent members of the species in vital interaction networks, retention of common properties across generations helps sustain the species. This side-effect of sexual reproduction may have also contributed to the very evolution and pervasiveness of sexual reproduction in nature.
@@ 1 Introduction
Natural reproduction is of course about transferring parental features to offspring. It is also well accepted that bi-parental reproduction yields resemblance of offspring in certain aspects to one parent or the other, or to both, or to having a mixture of traits from both parents. In this paper we explore a finer effect of sexual reproduction which is not addressed explicitly in discussions of evolution of sexual reproduction like those cited below, namely, that bi-parental reproduction conserves traits that are shared by both parents, and that may be common to the entire parent generation within the species.
While there are multiple theories for the sustainment and evolutionary effects of bi-parental reproduction, both sexual (with two or more mating types) and unisexual (with one mating type, as is common, for example in fungi [Heitman, 2015]), over mono-parental, clonal, asexual reproduction, the underlying question is still considered open [MacPherson et al., 2023, MacPherson et al., 2021, Livnat and Papadimitriou, 2016, Billiard et al., 2012]. and others.
The most commonly cited consequences of bi-parental reproduction include: (i) the introduction of diversity, which then leads to increased adaptivity to changing conditions, and enhanced competitiveness and the ability to co-evolve with pathogens, parasites, or symbionts; for example the Red Queen hypothesis [Van Valen, 1973]; (ii) the constraining the irreversible penetration of deleterious mutations, often termed Muller’s ratchet [Kondrashov, 1988, Muller, 1932]; (iii) preservation of common properties by facilitating the raising of reproduction barriers between divergent subgroups within a species [Barton et al., 2007, CH.22], DNA repair [Bernstein et al., 1981], [Barton et al., 2007, Ch.12 p.335] that is part of meiosis, sexual selection which promotes reproduction of individuals with certain preferred properties, and interference with specialization and speciation of variants that find and prefer new niches [Barton et al., 2007, Ch.16]. Indeed, the genome homogenization effect of bi-parental reproduction is often taken for granted when a population is panmictic, i.e., every organism of the species can mate with every other of the same species, and confined within a geographical region [Fitzpatrick et al., 2008, Abstract]; it is also often implicit from discussions of gene flow in structured populations [Barton et al., 2007, Ch.16 p.441.], or in pondering the puzzle of sympatric speciation, i.e., why do panmictic populations in a given location sometime give rise to diversity and speciation [Fitzpatrick et al., 2008, Mayr, 1982].
Here we show that in a population whose clonal reproduction is imperfect and may induce some change, as is the case with many kinds of DNA mutations [Barton et al., 2007, Ch.12], horizontal gene transfer [Keeling and Palmer, 2008] incorporation of viral genetic material [Koonin, 2010], etc., bi-parental reproduction that employs merging of encodings of inherited properties, has an underlying mathematical conservation effect that is independent of biological specifics, and is more general than the homogenization mechanisms listed above. Specifically, bi-parental reproduction conserves the properties that are common within the parents’ generation better than mono-parental reproduction. This, of course, then extends to conservation of ancestral properties across multiple generations. This view can be aligned with the observation in [Cohen, 2016] that one of the effects of sexual reproduction is “to maintain a frequency distribution of genes within the species that fits the life style and ecosystem”.
We support our claim that the effects of bi-parental reproduction are intrinsic, by initially showing that they exist in an abstracted setting: we deal with only a single property, and the mechanism does not attempt to model in detail meiosis, genetic recombination, epistasis, syngamy, the role of mating types, the role of the environment, and other complex aspects of reproduction in nature. While all these factors are of course important and should be part of realistic models of reproduction in nature, the particular aspects of bi-parental reproduction discussed here deserve a special focus beyond the published theories, mechanisms and effects.
1.1 Current theories do not imply conservation of shared traits.
Below we list some distinctions between offspring’s retention of traits that are common to their two parents, from presently published theories and observations around sexual reproduction.
@@
    Consider at first the prominent effect of sexual reproduction, which is the enhancing of species diversity. This capability relies on the fact that individuals differ from each other in the first place. By contrast, the effect discussed here focuses on the traits that do not change
@@
    Plain merging of encodings is not the same as repair: It does not involve a preferred template, and does not call for assessment of whether a structure is broken or intact.
@@
    Conservation of shared properties is not about elimination of traits that interfere with fitness. In the present model, entity survival is not affected by whether its traits are deleterious or beneficial to its ability to reproduce, or its rate of reproduction.
  4. 4.
    There is no preferential attitude to properties, in contrast, for example, to how sexual selection contributes to species homogenization.
  5. 5.
    The present analysis is different from the homogenization effect of reproductive barrier raised between divergent subspecies, as the group we analyze is the entire population where every individual can and may mate with every other individual in the species.
  6. 6.
    The way sexual reproduction and associated gene flow constrains specialization and fitting into new niches by diluting new beneficial traits may be a special case of the mechanism described here. Bi-parental reproduction slows down property drift regardless of how useful the new properties may be in new niches. We emphasize that interference with speciation comes from retention of shared traits at the expense of diverse ones. Hypothetically, the dilution of over-specialized properties could occur by further randomizing then, and preventing them from becoming common and shared, without slowing down the drifting away of the properties that are shared and common.
  7. 7.
    Most importantly, homogenization of the next generation as induced by sexual reproduction, and discussed extensively in the literature, does not necessarily imply that a shared property of a homogeneous new generation is also closer to the shared property of the immediately preceding parent generation. For example, if the parent generation is comprised of clusters of different trait values, a homogeneous child generation may still have genotype and phenotype properties that are not in any of these clusters.
@@
In the next section, we briefly discuss some of the motivation for the particular principles applied in the ensuing model. In Section 3 we use a synthetic model with encodings consisting of sets of real numbers to prove mathematically the inevitability of conservation of parental properties as a general effect of bi-parental reproduction derived from basic probability and the central limit theorem. In Section 5 we illustrate the effect with examples drawn from simulation runs that reproduce diverse objects including abstract sets of points, images of printed text, and images of flowers.
In summary, given that survival of organisms, species and entire ecosystems depends on organisms fitting in into a rich interaction network, and given that all living individuals are destined to eventually die, our interim results show that intrinsic properties of bi-parental reproduction likely help sustain each species by enabling new generations to carry out the interaction roles performed by their parents and by several preceding generations. Further studies are needed to generalize and expand the model, and to confirm that the conservation of shared traits by bi-parental reproduction is indeed a distinct natural phenomenon.
@@ 2.1 Codes, interaction codes, and autoencoding
The present work is part of a larger research project. In [Cohen and Marron, 2023, Cohen and Marron, 2020] and related papers we have outlined how sustainment of a species depends on interaction networks in which it participates. In such networks, the role of each species is naturally and automatically encoded in what we term species interaction code, manifested in the genes, physiology and environment of the individuals of the species. We aim to extend our modeling to address not only species properties, but inter- and intra-species interactions, and research how sexual reproduction may help encode and conserve species interaction codes over multiple generations.
As part of our work on natural autoencoding we have experimented with autoencoding of basic structures like images of straight line segments [Bayer et al., 2023], which hypothetically, can be encoded in just four numbers: the two pairs of coordinates of the endpoints, or a pair of coordinates, an angle and a length. While not dealing with autoencoding per-se, the present model further reduces this two dimensional line encoding problem to a one dimensional continuous range of real numbers which can be encoded in just two numbers, like the value of its boundaries, or its center point and its length. For the formal proof we used a normal distribution with a given mean and variance, rather than a continuous segment.
@@
In this spirit, we focus on the passing down of the encoding of a trait, not the trait itself, even though the phenotype is what in the end sustains the organism. For example, we assume that if the property of interest of an individual is its average height as an adult, or its expected life span, or its varying colors in various seasons, then this property must somehow be encoded before it can be recreated. The encoding itself does not have to be compact or minimalistic; for example, it may involve extensive redundancy, to overcome decoding errors, or may be captured physically, as is the case in a human-made diagram or graph, which may be recorded in many pixels.
@@
This perspective connects the present work to quantitative genetics [Barton et al., 2007, Ch.14;16;17;19;26],[Kang, 2020],[Fisher, 1918]. While blending inheritance theory, where parents traits are mixed to form the offspring features, is long considered obsolete, quantitative genetics studies the many quantitative phenotypic traits, like lifespan, weight, strength, or speed, which are affected by vast genetic information, commonly represented in Quantitative Trait Loci (QTL) which may include multiple genes, regulatory sequences, epigenetic effects, etc. The phenotype aggregation of such polygenic features involves among others operations that resemble adding and averaging the various genetic factors within the organism. Furthermore, the way offspring QTL information is determined from the merging of parents’ QTL is known to be highly complex [Cui et al., 2004, Niu et al., 2022], and cannot yet be reduced to a small set of rules; the model makes a modest attempt to reflect this complexity.
@@
Throughout this paper, we use the term an encoding to refer to an object or structure that encodes, or represents, another object or concept. Thus, an encoding is synonymous here with the term a code, which we presently reserve for the broader concept of species interaction code.
@@
While the model does not aim at mimicking meiosis and creation of gametes, the model includes an elaborate step of recreating a parent’s encoding that will then serve as a seed for subsequent reproduction steps; this is instead of just directly copying the parent’s encoding.
@@ 2.2 Information aggregation by set operations
Note that pure averaging of a single property represented by a random variable would be analogous to combining two one-dimensional random walks, with the same step size and equal probabilities for stepping in either direction. Such random walks can be readily shown to have a mean position which is the average of the two origins, and, more importantly, its variance is half the (equal) variance of the two original random walks. This is aligned with the known fact that bi-parental reproduction reduces the variance within the offspring generation.
@@ 2.3 Minimalism, abstraction and accessible language
The model is not intended to be a full model of what happens in nature. In fact, we try to show that in situations of noisy inheritance, it is the very merging of parent encodings that enables the cross-generation conservation of ancestral traits. Examples of elements of the abstraction that we employed include: (i) The property of interest and its encoding both involve real numbers. Extending the arguments to pure symbolic entities, or concrete objects like molecules, is left for a future research; (ii) The inevitable variation in natural reproduction is modeled by replacing direct copying with sampling from a normal distribution; (iii) the bi-parental reproduction is unisexual, i.e., there are no distinct mating types.
Since here we aim to only show relations between certain values, rather than establishing those values, we adopt the “story proof” argumentation style, as defined and used in[Blitzstein and Hwang, 2015]; this style relies on text of intuitive, yet well founded, logic statements to support the claims, and keeps mathematical formulas to a minimum. In addition, the occasional use of terms that are commonly applied to populations of living organisms is only for convenience.
@@ 3.1 The formal model
Consider the following synthetic model of a population of entities, also referred to as individuals. The size of the population is kept fixed at a number $n$. The analysis does not include considerations of time. The population reproduces in methods discussed below. A set of $n$ entities, referred to as the cohort of the next, or child, generation, appears instantly and synchronously.
@@
We label the individual entities in a cohort as $\{e_{1},e_{2},...,e_{n}\}$;
All entities are associated with exactly one “phenotype” property $P$, which is manifested as a real number representing some quantity. Different entities may have the same or different values for $P$. For illustration, the abstract entities may be thought of as fixed-width sticks where $P$ is their length, or as lumps of material where $P$ is their weight. Let $p$ denote any value of $P$, and, for example, $p_{i}$ is the value of $P$ in the entity $e_{i}$. As for the focus on a quantitative property, see [Barton et al., 2007, Ch.16 p.441]. With each new generation, under the two reproduction methods, the $p_{i}$ values and their distribution within each new cohort may change. In each cohort $c$ we are first interested in the mean of the $p_{i}$ values within the cohort
$p_{c}=\frac{1}{n}\sum_{i=1}^{n}p_{i}~{}~{}.$
In each parent generation $c$, one can view $p_{c}$ as representing a common property, or capability, which is manifested with some diversity in the living individuals. And, if the value of $p$ in an individual is essential for survival within the species’ ecosystem, then the distribution of the property $P$ among the members of the next generation $c^{\prime}$ affects the likelihood that individuals from $c^{\prime}$ will be able to participate in ecological networks in which their parents from $c$ no longer function.
The “genotype” encoding of $P$ in each individual $e_{i}$ is a set $g_{i}$ of $k$ of real values which are also referred to as the points in $g_{i}$. We allow the rare case of any two points in a set being equal to each other, without resorting to other mathematical terms like $g_{i}$ being a multiset or a tuple. The integer $k$ is a fixed parameter in the model, and it is large, say, $k\geq 100$; we also constrain $k$ to be an even number.
The phenotype manifestation of the property $P$ in $e_{i}$ is the mean of the points in $g_{i}$. That is, given an individual $e_{i}$, $1\leq i\leq n$, with genotype encoding $g_{i}=\{x_{i_{1}},x_{i_{2}},...,x_{i_{k}}\}$,its phenotype value $p_{i}$ is computed as follows:
$p_{i}=\frac{1}{k}\sum_{\ell=1}^{k}x_{i_{\ell}}~{}~{}.$

### v5

# Natural Averaging May Complement Known Biological Constraints in Bi-parental Reproduction’s Advantages Over Mono-parental in Conserving Species Quantitative Traits

@@ Abstract
Commonly recognized evolutionarily relevant effects of sexual reproduction include increased diversity, accelerated adaptation, and constrained accumulation of deleterious mutations, along with a secondary effect of species genotype homogenization. Still, strong published arguments prioritize the contribution of biological mechanisms underlying bi-parental reproduction to maintaining species identity above their contribution to diversity. Here, we contribute to the latter position. In an initial mathematical analysis and simulation, we show that in an environment where copying is prone to error, quantitative polygenic traits that are shared within a parents’ generation are transmitted to future generations under bi-parental reproduction with less deviation than under asexual reproduction. Furthermore, we abstract away many biological details, and show that this trait conservation is a general statistical effect, driven by the very nature of mixing of parental traits, separately from DNA repair and from the reproductive failures, barriers and disadvantages induced by biological mechanisms. Since survival of ecosystem interaction networks depends on the ability of individuals to replace the networked function of failing, dying or absent members of the same species, more faithful inheritance of common traits helps sustain species and ecosystems. This sustaining effect may have contributed to the very evolution of sexual reproduction.
keywords:
Sexual reproduction, Asexual reproduction, Bi-parental, Uniparental, Mono-parental, Unisexual, Clonal, Autoencoding, Probabilistic Sampling
@@ 1 Introduction
There are multiple theories for the evolution of bi-parental reproduction, both sexual (with two or more mating types) and unisexual (with one mating type, as is common, for example in fungi [1]), as contrasted with mono-parental, clonal, asexual reproduction. However, the underlying question is still considered open [2, 3, 4, 5]. Presently, in scientific publications and popular discussions covering the drivers for evolution of sexual reproduction, or the consequences of various types of bi-parental reproduction, the most commonly listed aspects are:
@@
    Introduction of diversity, which then leads to increased adaptivity to changing conditions, enhanced competitiveness and the ability to co-evolve with pathogens, parasites, or symbionts; for example, the Red Queen hypothesis [6];
@@
    Constraining of the irreversible penetration of deleterious mutations, often termed Muller’s ratchet [7]; this occurs by combination of dominant/recessive relations of pairs of alleles, and by the reproductive disadvantage imposed by the deleterious trait at hand, which may be augmented by Kondrashov’s Hatchet[8] – interactions between slightly deleterious mutations; and,
@@
    Preservation of common traits through biological processes including (i) the raising of reproduction barriers between divergent subgroups within a species [9], (ii) DNA repair that occurs as part of meiosis [10, 9], (iii) sexual selection which promotes reproduction of individuals with certain preferred traits[9], and (iv) interference with specialization and speciation of variants that find and prefer new niches [9](ch.16).
The relative importance or prominence of the various factors is under debate. For example, in [11], Heng and Gorelick first observe that diversity, adaptivity, and constraining of deleterious mutations are generally considered the more prominent factors, while the contribution of sexual reproduction to faithful copying and genotype homogenization is relegated to a secondary or supplementary role; they then argue that much greater significance should be assigned to the latter, as these processes maintain the species identity – the functions that constitute the delicate interaction networks that sustain the species and its ecosystem. This view can be aligned with the observation in [12] that one of the effects of sexual reproduction is “to maintain a frequency distribution of genes within the species that fits the life style and ecosystem”.
In this paper we contribute to this emphasis on retention of common parent-generation traits (hereafter abbreviated as Retention of Common Traits (RoCT)) with generalized statistical analysis and simulation. We show that in a population where any kind of copying is imperfect, bi-parental reproduction that includes merging of encodings of inherited traits, conserves in the offspring generation traits that are common within the parents’ generation better than does mono-parental reproduction. Furthermore, this effect is a general statistical property, driven by the very nature of mixing of parental traits, and distinct from the biological processes of repair and limits on reproduction. This, of course, then extends to conservation of traits across multiple generations. Given that survival of organisms, species and ecosystems depends on interaction networks, and that all organisms eventually die, higher fidelity RoCT helps sustain species by enabling offspring to step in and carry out interaction roles performed by their ancestors.
While mathematical/statistical comparisons of RoCT under mono- and bi-parental reproduction, separately from the particular biological mechanisms, are quite sparse in recent literature, certain aspects of the topic were discussed as far back as a century ago. For example, in 1918, Fisher [13] establishes the correlation between the phenotypic expression of quantitative traits (also termed characters) in offspring and that of the parents. In 1886, Galton [14] reports observations that for quantitative traits, offspring traits tended towards the mean of the population, eliminating extremes that may have been manifested by their parents, a phenomenon termed “regression towards the mean”. In his experiment measuring human stature he also observes population-wide processes which he summarizes as “two opposite sets of actions, one concentrative and the other dispersive", but concludes that “they necessarily neutralise one another, and fall into a state of stable equilibrium”[14](p.256).
Hamilton [15], observes that “sex reduces the fitness variance in frequency-dependent host-parasite systems”; this may appear relevant to the present discussion if fitness is equated with the trait at hand. However, this model is unique in that the fitness variance is reduced in an environment where a parasite evolves to attack the phenotype with highest frequency, because sexual reproduction creates a population with diverse traits, enabling evasion of the parasite’s attacks.
Doebeli [16] compares mathematical models for dynamics of population size, i.e., density fluctuations, under various parameters including population diversity, and under sexual and asexual reproduction. The focus is on quantitative phenotypic traits, which are manifested as fitness. The entire setting is that of competition and natural selection, with models addressing competition between individuals in the population and between sexuals and asexuals within the population. The model also considers population regulation through host-parasite interactions. The models and simulations show that sexual reproduction reduces density fluctuations and fitness variance within a generation. One may also infer that the fitness of an offspring generation is closer to that of the parent generation under sexual reproduction than under asexual reproduction, which alludes in some ways to RoCT.
Such analyses focus either on statistical analysis of trait inheritance in the absence of mutation, or on the evolution of traits when both genetic changes and natural selection, or fitness, play a central role. The present paper presents an analytical and simulation perspective on inherent advantages of sexual reproduction over asexual in preserving common traits when the copying mechanisms are imperfect, say, due to mutations, but in the absence of biological considerations, i.e., not considering repair mechanisms or whether the trait at hand contributes to or interferes with reproductive success.
RoCT may be correlated with homogenization of offspring genotype, but these are distinct concepts. Indeed, the fact that bi-parental reproduction results in homogenization of the genome in the offspring generation is often taken for granted when a population is panmictic, i.e., every organism of the species can mate with every other of the same species, and all are confined within a geographical region (see, e.g., the abstract of [17]); it is also often implicit from discussions of gene flow in structured populations (see, e.g., ch.16, p.441 in [9]), or in pondering the puzzle of sympatric speciation: why panmictic populations sometimes give rise to diversity and speciation even within a given location [17, 18]. However, mere homogenization within the offspring generation alone does not immediately imply retention of parental traits and may even dilute such traits.
As in the research by Doebeli [16], our analysis and simulation is constrained to quantitative traits, which by and large are polygenic. Quantitative traits constitute the majority of the identifiable traits in organisms [19](p.239), and our knowledge about how various genes affect each such phenotypic trait is very limited. Indeed, the very existence of quantitative phenotypic traits whose manifestation can take into account summing or averaging the traits of the two parents is an important factor that enables the retention of what is common or shared in the entire generation of these parents.
@@
The paper is structured as follows: In Section 2, we discuss some of the motivation for the particular principles applied in the ensuing modeling, analysis and simulation examples. In Section 3 we introduce a model for genotype, phenotype and sexual and asexual transmittal of quantitative characters. In Section 4 we prove, using basic probability and the central limit theorem in a restricted synthetic model, the inevitability of conservation of quantitative parental traits as a general effect of bi-parental reproduction, that is independent of biological specifics, and is more general than known genotype homogenization mechanisms. In Section 5 we illustrate the differences between the RoCT effect under bi- and mono-parental reproduction with examples drawn from simulation runs that reproduce diverse objects including abstract sets of points, images of printed text, and images of flowers. The latter simulations also illustrate how conservation of many distinct traits, like individual Red-Green-Blue (RGB) values in thousands of pixels, can affect conservation of emergent traits, like recognizing that an image is of a flower.
Further studies are needed to generalize and expand the model, and to confirm that the improved RoCT afforded by bi-parental reproduction is indeed a distinct natural phenomenon.
@@ 2 Modeling Rationale
We support our claim that the effects of bi-parental reproduction are intrinsic, by initially showing that they exist in an abstracted setting: we deal with only a single quantitative trait; we do not attempt to model in detail transmittal of non-quantitative traits, or the processes of meiosis, genetic recombination, epistasis, and syngamy, the role of mating types, the role of the environment, or other complex aspects of reproduction in nature. While all these factors are important and should be part of realistic models of reproduction in nature, the particular aspects of bi-parental reproduction discussed here deserve a special focus beyond the published theories, mechanisms, and observed effects.
@@ 2.1 Codes, interaction codes, and autoencoding
The present work is part of a larger research project. In [20, 21] and related papers we have outlined how sustainment of a species depends on interaction networks in which it participates. In such networks, the role of each species is naturally and automatically encoded in what we term species interaction code, manifested in the genes, physiology and environment of the individuals of the species. As interactions depend on traits that enable them, faithful transmittal is key.
@@
An important assumption in the analysis is that to be transmitted, a phenotypic trait, like height, life span, or coloring, must first be encoded. The encoding itself does not have to be compact or minimalistic; for example, it may involve extensive redundancy in order to overcome decoding errors, or it may be captured physically, as is the case in a human-made diagram, or graph, of a simple relationship between a few objects, which may be physically recorded in millions of pixels.
@@
This perspective connects the present work to quantitative genetics [9] (ch.14,16,17,19,26),[22, 13]. While blending inheritance theory, where parents traits are mixed to form the offspring features, is long considered obsolete when discussing individual genes and alleles  [23](p.52), quantitative genetics studies the many quantitative phenotypic traits which may manifest averaging effects. Such traits include lifespan, weight, strength, speed, and many others. Quantitative traits are affected by vast genetic information, commonly represented in Quantitative Trait Loci (QTL) which may include multiple genes, regulatory sequences, epigenetic effects, etc. The phenotype polygenic aggregation involves, among others, operations that resemble adding and averaging the various genetic factors within the organism. Furthermore, the way offspring QTL information is determined from the merging of parents’ QTL is known to be highly complex [24, 25], and cannot yet be reduced to a small set of rules; the model makes a modest attempt to reflect this complexity.
@@
Throughout this paper, we use the term an encoding to refer to an object or structure that encodes, or represents, another object or concept. Thus, an encoding is synonymous here with the term a code, which we presently reserve for the broader concept of species interaction code. The use of the term "an encoding" as an instance, distinguishes it from the concept of “a code” as is used by Barbieri and others, where it refers to the entire mapping between domains [26].
@@
While the model does not aim to mimic meiosis or gamete formation, we opted to include in the model an elaborate step of recreating a parent’s encoding that will then serve as a seed for subsequent reproduction steps, rather than sufficing with straightforward copying of the parent’s encoding.
@@ 2.2 Information aggregation by set operations
Note that pure averaging of a single trait represented by a random variable whose copying is subject to error with some probability, would be analogous to combining two one-dimensional random walks, with the same step size and equal probabilities for stepping in either direction. When two such random walks, starting at different origins, are combined by stepping in parallel and averaging the location, the resulting mean position is the average of the two origins. More importantly, the position variance is half the (equal) variance of the two original random walks. This is aligned with the known fact that bi-parental reproduction reduces the variance within the offspring generation.
@@ 2.3 Minimalism, abstraction and accessible language
The model is not intended to be a full model of what happens in nature. In fact, we try to show that in situations of noisy inheritance, it is the very merging of parent encodings that enables the cross-generation conservation of ancestral traits. Examples of elements of the abstraction that we employed include: (i) The trait of interest and its encoding both involve real numbers. Extending the arguments to pure symbolic entities, or concrete objects like molecules, is left for a future research; (ii) The inevitable variation in natural reproduction is modeled by replacing direct copying with sampling from a normal distribution around the source value; (iii) the bi-parental reproduction is unisexual, i.e., there are no distinct mating types.
Since here we aim to only show relations between certain values, rather than establishing those values, we adopt the “story proof” argumentation style, as defined and used in [27]; this style relies on text of intuitive, yet well founded, logic statements to support the claims, and keeps mathematical formulas to a minimum. In addition, the occasional use of terms that are commonly applied to populations of living organisms is only for convenience.
In an attempt to align with natural processes, in the mathematical analysis and the simulation examples below, the summing and averaging of encoded traits is not carried out by of the very process of of fusing parental genetic material (syngamy, in nature), but by the organism development process that interprets the combined codes. Furthermore, the encoding that supports this process works well also for mono-parental reproduction.
@@ 3.1 The formal model
Consider the following synthetic model of a population of entities, also referred to as individuals. The size of the population is kept fixed at a number $n$. We label the entities in a cohort as $\{e_{1},e_{2},...,e_{n}\}$. The analysis does not include considerations of time. The population reproduces in methods discussed in detail below; most basically, in every reproduction step, a set of $n$ entities, referred to as the cohort of the next, or child, generation, appears instantly and synchronously.
@@
All entities are associated with exactly one “phenotype” quantitative trait $P$, which is manifested as a real number. Different entities may have the same or different values for $P$. For illustration, the abstract entities may be thought of as fixed-width sticks where $P$ is their length, or as lumps of material where $P$ is their weight. Let $p$ denote any value of $P$, and, for example, $p_{i}$ is the value of $P$ in the entity $e_{i}$. With each new generation, under the two reproduction methods, the $p_{i}$ values and their distribution within each new cohort may change. In each cohort $c$ we are first interested in the mean $p_{c}$ of the $p_{i}$ values within the cohort:
$p_{c}=\frac{1}{n}\sum_{i=1}^{n}p_{i}\leavevmode\nobreak\ \leavevmode\nobreak\ .$
In each parent generation $c$, one can view $p_{c}$ as representing a common trait, or capability, which is manifested with some diversity in the living individuals. (If the value of $p$ in an individual is essential for survival within the species’ ecosystem, then the distribution of the trait $P$ among the members of the next generation $c^{\prime}$ affects the likelihood that individuals from $c^{\prime}$ will be able to participate in ecological networks in which their parents from $c$ no longer function. Note however, that such survival considerations are not part of the model, analysis, and simulations.)
The “genotype” encoding of $P$ in each individual $e_{i}$ is a set $g_{i}$ of $k$ real values which are also referred to as the points in $g_{i}$. We allow the rare case of any two points in a set being equal to each other, without resorting to other mathematical terms like $g_{i}$ being a multiset or a tuple. The integer $k$ is a fixed parameter in the analytical model, and it is large, say, $k\geq 100$; we also constrain $k$ to be an even number.
The phenotype manifestation of the trait $P$ in $e_{i}$ is the mean of the points in $g_{i}$. That is, given an individual $e_{i}$, $1\leq i\leq n$, with genotype encoding $g_{i}=\{x_{i_{1}},x_{i_{2}},...,x_{i_{k}}\}$,its phenotype value $p_{i}$ is computed as follows:
$p_{i}=\frac{1}{k}\sum_{\ell=1}^{k}x_{i_{\ell}}\leavevmode\nobreak\ \leavevmode\nobreak\ .$